Find a polynomial equation of lowest degree with integer coefficients and has 0, 3, and plus/minus sq rt of 2 as zeros. Please, anyone help.
Answer by venugopalramana(3286)

Find a polynomial equation of lowest degree with integer coefficients and has 0, 3, and plus/minus sq rt of 2 as zeros. Please, anyone help.
ROOTS ARE ...0,3,SQRT(2),-SQRT(2)...SO EQN IS
(X-0)(X-3)(X-SQRT2)(X+SQRT2)=0
(X^2-3X)(X^2-2)=0
X^4-3X^3-2X^2+6X=0